Question Unit 6 Tutorials If you need hekp Question 4 O Mark this question Joan conducted a study to see how common binge drinking is on her college campus She defined frequent binge drinking as having five or more drinks in a row three or more times in the past two ents who replied to her survey, 64 fit this criterion. Joan wants to construct à significance test for her data She finds that the proportion of binge drinkers nationally is 13.19% The z statistic for this data is z= pa
Answer #1

Given that, n = 593 and x = 64

sample proportion (hat p) = x/n = 64/593 = 0.1079

population proportion ( p ) = 13.1% = 0.131

=> q = 1 - p = 1 - 0.131 = 0.869

The z test statistic for this data is,

Z = p-p-0.1079-0. 131 0.1310,869

Z = -1.67